# Nonequilibrium Probes of Quantum Geometry in Gapless Systems

### v1

@@ Appendix E Lattice calculations
$\Bigl|\langle\psi_{N}(0)|\psi_{N}(T)\rangle\Bigr|^{2}=\Bigl|\langle\psi_{N}(0)|U_{N}(T)|\psi_{N}(0)\rangle\Bigr|^{2},\qquad U_{N}(T)\equiv\overset{\longleftarrow}{{\cal T}}e^{-i\int_{0}^{T}\textrm{d}t\,H_{N}(t)}$ (E.2)
after one period $T$. To this end, as is commonplace, we exploit the fact that the Hamiltonian for any periodic fermionic chain with nearest-neighbor couplings can be represented by a periodic Jacobi matrix. In our case, since there is no external field, this appears as the adjacency matrix
$\mathsf{J}(t)\equiv\begin{pmatrix}0&J_{1}(t)&0&\ldots&0&J_{N}(t)\\ J_{1}(t)&0&J_{2}(t)&\ldots&0&0\\ 0&J_{2}(t)&0&\ldots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&0&J_{N-1}(t)\\ J_{N}(t)&0&0&\ldots&J_{N-1}(t)&0\\ \end{pmatrix}$, (E.3)
which allows us to write $H_{N}(t)=(c_{1}^{\dagger},c_{2}^{\dagger},\ldots,c_{N}^{\dagger})\mathsf{H}(t)(c_{1}^{\vphantom{\dagger}},c_{2}^{\vphantom{\dagger}},\ldots c_{N}^{\vphantom{\dagger}})^{T}$ with $\mathsf{H}(t)\equiv-\mathsf{J}(t)/2$ interpreted as a single-particle Hamiltonian. Let $\mathsf{U}$ be the unitary matrix that diagonalizes $\mathsf{H}(t)$ at time $t=0$. It follows that
$H_{N}(0)=\sum_{j=1}^{N}E_{j}(0)\gamma_{j}^{\dagger}\gamma_{j}^{\vphantom{\dagger}},\quad\gamma_{j}=\sum_{\ell=1}^{N}(\mathsf{U}^{\dagger})_{j\ell}c_{\ell}$ (E.4)
with eigenenergies $E_{j}(0)$ of $\mathsf{H}(0)$ and new fermionic creation (annihilation) operators $\gamma_{j}^{\dagger}$ ($\gamma_{j}^{\vphantom{\dagger}}$). General properties for periodic Jacobi matrices [75] imply that the $E_{j}(0)$ can be ordered to be (nonstrictly) increasing in $j=1,\ldots,N$ (with each eigenenergy equal to at most one other).
Recall that half filling corresponds to filling all single-particle states with negative energies. For simplicity, let $N$ be even. Then, all states with $j=1,\ldots,N/2$ have nonpositive energies, [^13] For $N$ even, let $S=\operatorname{diag}(1,-1,1,-1,\ldots,1,-1)$. Since $\mathsf{J}(t)$ in Eq. (E.3) satisfies $\mathsf{S}\mathsf{J}(t)=-\mathsf{J}(t)\mathsf{S}$, if $\mathbf{x}$ is an eigenvector of $\mathsf{H}(t)=-\mathsf{J}(t)/2$ with eigenvalue $E$, then $\mathsf{S}\mathbf{x}$ is also an eigenvector with eigenvalue E-E. i.e. the ground state at half-filling can be taken as $|\psi_{N}(0)\rangle=\prod_{j=1}^{N/2}\gamma_{j}^{\dagger}|\mathrm{vac}\rangle$, where $|\mathrm{vac}\rangle$ denotes the state with no fermions. (For $N$ odd, one must replace $N/2$ by the maximal value $N_{\text{max}}$ so that $E_{N_{\text{max}}}(0)$ is negative.) It follows that the return probability (E.2) reduces to computing
@@
$\Bigl|\langle\mathrm{vac}|\gamma_{1}^{\vphantom{\dagger}}\ldots\gamma_{N/2}^{\vphantom{\dagger}}U_{N}(T)\gamma_{N/2}^{\dagger}\ldots\gamma_{1}^{\dagger}|\mathrm{vac}\rangle\Bigr|^{2}\\ =\Bigl|\langle\mathrm{vac}|[U_{N}(T)^{\dagger}\gamma_{1}^{\vphantom{\dagger}}U_{N}(T)]\ldots[U_{N}(T)^{\dagger}\gamma_{N/2}^{\vphantom{\dagger}}U_{N}(T)]\gamma_{N/2}^{\dagger}\ldots\gamma_{1}^{\dagger}|\mathrm{vac}\rangle\Bigr|^{2}.$ (E.5)
@@
To find $U_{N}(T)^{\dagger}\gamma_{j}U_{N}(T)$, we need to discretize the continuous time evolution: we approximate the evolution under $H_{N}(t)$ for one period as one consisting of $M\gg 1$ time steps, each evolved under a static Hamiltonian $H_{N}(m\delta t)$ for a time $\delta t=T/M$. In other words,
$U_{N}(T)\approx\prod_{m=0}^{M}e^{-iH_{N}(m\delta t)\delta t}.$ (E.6)
@@
As when $t=0$, the single-particle Hamiltonian $\mathsf{H}(m\delta t)$ corresponding to $H_{N}(m\delta t)$ can be diagonalized by a unitary matrix $\mathsf{V}_{m}$. Let $E_{j}(m\delta t)$ denote the eigenenergies of $\mathsf{H}(m\delta t)$. Defining $\mathsf{W}_{m}(\delta t)\equiv\mathsf{U}^{\dagger}\mathsf{V}_{m}^{\vphantom{\dagger}}\operatorname{diag}\bigl(e^{-iE_{1}(m\delta t)\delta t},\ldots,e^{-iE_{N}(m\delta t)\delta t}\bigr)\mathsf{V}_{m}^{\dagger}\mathsf{U}$, it follows that
$e^{iH_{N}(m\delta t)\delta t}\gamma_{j}e^{-iH_{N}(m\delta t)\delta t}=\sum_{\ell=1}^{N}\Bigr(\mathsf{W}_{m}(\delta t)\Bigr)_{j\ell}\gamma_{\ell}.$ (E.7)
@@
We use this decomposition for each time step in Eq. (E.5), which reduces the whole calculation to evaluating a correlator for the $\gamma_{\ell}$s. The return probability can then be approximated as
@@
$\Bigl|\langle\psi_{N}(0)|\psi_{N}(T)\rangle\Bigr|^{2}\approx\left|\det_{1\leq i,\ell\leq N/2}\left(\prod_{m=0}^{M}\mathsf{W}_{m}(\delta t)\right)_{j\ell}\right|^{2}$ (E.8)
with the help of Wick’s theorem.

### v2

@@ Appendix E Lattice calculations
$\begin{gathered}\Bigl|\langle\psi_{N}(0)|\psi_{N}(T)\rangle\Bigr|^{2}=\Bigl|\langle\psi_{N}(0)|U_{N}(T)|\psi_{N}(0)\rangle\Bigr|^{2},\\ U_{N}(T)\equiv\overset{\longleftarrow}{{\cal T}}\exp\!\Bigl(-i\int_{0}^{T}\textrm{d}t\,H_{N}(t)\Bigr).\end{gathered}$ (126)
In terms of the creation and annihilation operators introduced in Eq. (124), this can be recast as
$\Bigl|\langle\psi_{N}(0)|\psi_{N}(T)\rangle\Bigr|^{2}$
@@
$=\Bigl|\langle\mathrm{vac}|\gamma_{1}^{\vphantom{\dagger}}\ldots\gamma_{N/2}^{\vphantom{\dagger}}U_{N}(T)\gamma_{N/2}^{\dagger}\ldots\gamma_{1}^{\dagger}|\mathrm{vac}\rangle\Bigr|^{2}$
$=\Bigl|\langle\mathrm{vac}|[U_{N}(T)^{\dagger}\gamma_{1}^{\vphantom{\dagger}}U_{N}(T)]\ldots[U_{N}(T)^{\dagger}\gamma_{N/2}^{\vphantom{\dagger}}U_{N}(T)]$
$\quad\;\times\gamma_{N/2}^{\dagger}\ldots\gamma_{1}^{\dagger}|\mathrm{vac}\rangle\Bigr|^{2},$ (127)
@@
since $U_{N}(T)$ acting on $\langle\mathrm{vac}|$ only gives a phase. To compute each factor $U_{N}(T)^{\dagger}\gamma_{j}U_{N}(T)$ here, one needs to discretize the continuous time evolution. This is done as in Eq. (75) by approximating the evolution under $H_{N}(t)$ as consisting of $M\gg 1$ time steps, each evolved under a static Hamiltonian $H_{N}(m\delta t)$ for a time $\delta t=T/M$. In other words,
$U_{N}(T)\approx\prod_{m=0}^{M}e^{-iH_{N}(m\delta t)\delta t},$ (128)
where the approximation improves as $M$ increases.
@@
Let $\mathsf{V}_{m}$ be the unitary $N\times N$ matrix that diagonalizes the single-particle matrix Hamiltonian $\mathsf{H}(m\delta t)$ corresponding to $H_{N}(m\delta t)$. Let also $E_{j}(m\delta t)$ denote the eigenenergies of $\mathsf{H}(m\delta t)$. Defining $\mathsf{W}_{m}(\delta t)\equiv\mathsf{U}^{\dagger}\mathsf{V}_{m}^{\vphantom{\dagger}}\operatorname{diag}\bigl(e^{-iE_{1}(m\delta t)\delta t},\ldots,e^{-iE_{N}(m\delta t)\delta t}\bigr)\mathsf{V}_{m}^{\dagger}\mathsf{U}$, it follows that
$e^{iH_{N}(m\delta t)\delta t}\gamma_{j}e^{-iH_{N}(m\delta t)\delta t}=\sum_{\ell=1}^{N}\Bigr(\mathsf{W}_{m}(\delta t)\Bigr)_{\!j\ell\,}\gamma_{\ell}.$ (129)
@@
We use this decomposition for each time step in Eq. (127), which reduces the whole calculation to evaluating a correlator of the $\gamma_{\ell}$s. The return probability (127) can finally be approximated as
@@
$\Bigl|\langle\psi_{N}(0)|\psi_{N}(T)\rangle\Bigr|^{2}\approx\left|\det_{1\leq i,\ell\leq N/2}\left(\prod_{m=0}^{M}\mathsf{W}_{m}(\delta t)\right)_{j\ell}\right|^{2}$ (130)
with the help of Wick’s theorem. This is the approximate formula used to plot the numerical lattice results in Figs. 1 and 6.
To relate the above to our CFT results, let $J_{j}(t)=[J_{t}(x_{j})+J_{t}(x_{j+1})]/2$ be given by a slowly varying smooth function $J_{t}(x)$ of $x$, for each time $t$. The hopping (or coupling) between sites $j$ and $j+1$ is thus attributed to the sites themselves. The CFT propagation velocity, emerging when linearizing the dispersion relation around its two Fermi points, can then be expressed as $v_{t}(x_{j})=J_{t}(x_{j})(L/N)\sin(k_{F}L/N)$, where $k_{F}$ is the Fermi momentum and $L/N$ is the lattice spacing (cf. Ch. 5 in [107]). Exactly at half filling, these satisfy $k_{F}L/N=\pi/2$, in which case $v_{t}(x_{j})=J_{t}(x_{j})L/N$, as used in Figs. 1 and 6. That said, the same procedure is equally valid away from (but sufficiently near) half filling.